2- Consider a system containing pure hydrogen sulfide at 300 K and 20 bar. The following equation of state characterizes the PvT behavior of H2S well under these conditions: 0.422 1.6 PTc Pv RT 1 0.083 PCT T1.6 Using this equation of state, find the fugacity and fugacity coefficient.
Answer #1

イhe equation of state for drgen &ulfde is- 1 6 Tc= Critical temperature of H269-3733K tc= 11 pressure. 11 ニ89.70 bar = gg,

p2=20 bar =Qo xo-q87)ahu. = 19-74 at n P, = Wrut wary loo pressure C9o) ~o-1 atm 2. av inf 2.234 a e2839 Figaert coetf 2 ba C